OpenStudy (anonymous):

integration e^x/e^x+1 dx

OpenStudy (anonymous):

\[\Large \int \frac{e^x}{e^x+1}dx \] That? If yes, then use the u Substitution method for \[\Large u=e^x+1 \]

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{e^{x}}{e ^{x}+1 }\] but u=\[e ^{x}\]

OpenStudy (anonymous):

Easier to let u = e^x + 1

OpenStudy (anonymous):

You don't want to do a u=e^x substitution, this turns out to be redundant, so as I suggested u=e^x+1 will do you better.

OpenStudy (anonymous):

okay.. i will try u=e^x+1

OpenStudy (anonymous):

\[\Huge \checkmark \]

OpenStudy (anonymous):

then if \[x ^{3} \sqrt{x ^{2}+3}\] which is u please explain to me

OpenStudy (anonymous):

x^3 (x^2+3)^1/2

OpenStudy (anonymous):

\[\Large \int x^3\sqrt{x^2+3}dx \] If \[ \Large u=x^2+3 \] Then \[\Large x=\sqrt{u-3} \] So best might be using various substitutions: \[\Large dx=\frac{1}{2\sqrt{u-3}}du \] Maybe this will help you since the integral then becomes: \[\Large \int\sqrt{u-3}^3 \cdot \sqrt{u} \cdot \frac{1}{2\sqrt{u-3}}du \]

OpenStudy (anonymous):

If I didn't make any careless mistakes, then this will simplify to: \[\Large \frac{1}{2} \int (u-2)\cdot u^{1/2}du \]

OpenStudy (anonymous):

Sorry should be u-3

OpenStudy (anonymous):

Strange substitution at first, if you have any questions mention it to me @arifadli8

OpenStudy (anonymous):